OpenStudy (anonymous):

convert decimal to fraction .333= ?

OpenStudy (anonymous):

333/1000

OpenStudy (anonymous):

answer is 1/3 how??

OpenStudy (anonymous):

well 1/3 is .333..... that is .3 recurring

OpenStudy (anonymous):

Let x=0.3333..... 10x = 3.333....... Subtracting 2 from 1 9x = 3 x = 1/3

OpenStudy (anonymous):

0.333 is not the same as 0.333.....

OpenStudy (anonymous):

That is supposed to be a recurring.

OpenStudy (anonymous):

no amogh

OpenStudy (anonymous):

What i mean is, the question is wrong, its 0.3333....

OpenStudy (anonymous):

question is same and its here but wanted sm different view http://www.mathsisfun.com/converting-decimals-fractions.html see in last

OpenStudy (anonymous):

Yes, they have first mentioned a non-recurring number, 0.333 = 333/1000 [it is stated, it cannot be further simplified] And the second case is a recurring, 0.333....

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so amog ur answer is correct it seems .. this approch? Let x=0.3333..... 10x = 3.333....... Subtracting 2 from 1 9x = 3 x = 1/3

OpenStudy (anonymous):

amog's answer is correct for 0.333.... but .333 is different fraction

OpenStudy (anonymous):

0.3333..... goes on forever its an infinite decimal. sometimes it is written as 0.3 with a dot above the 3, but I cant do that on my keyboard!

OpenStudy (anonymous):

yup ....i got

OpenStudy (anonymous):

good - no probs